mansi is using an interactive map on her computer that allows her to zoom in or out. each time she zooms out the scale of the map increases by a power of ten. if she zooms out four times the scale is 10 to the 4th power times greater. write this number in standard form.
Answer by Theo(13342)

10^0 = 1
10^1 = 10
10^2 = 100
10^3 = 1,000
10^4 = 10,000

10^4 times greater is 10,000 times greater.